# Equal contribution

### v1

@@ Abstract
This paper presents a novel paradigm for building scalable 3D generative models utilizing pre-trained video diffusion models. The primary obstacle in developing foundation 3D generative models is the limited availability of 3D data. Unlike images, texts, or videos, 3D data are not readily accessible and are difficult to acquire. This results in a significant disparity in scale compared to the vast quantities of other types of data. To address this issue, we propose using a video diffusion model, trained with extensive volumes of text, images, and videos, as a knowledge source for 3D data. By unlocking its multi-view generative capabilities through fine-tuning, we generate a large-scale synthetic multi-view dataset to train a feed-forward 3D generative model. The proposed model, VFusion3D, trained on nearly 3M synthetic multi-view data, can generate a 3D asset from a single image in seconds and achieves superior performance when compared to current SOTA feed-forward 3D generative models, with users preferring our results over $70\%$ of the time.
@@ 3.2 Video Diffusion Models as 3D Multi-view Data Generator
The generated multi-view videos follow a similar camera trajectory to cover a ${}^{\circ}$ azimuth range, but the elevation depends on the generated condition image, which is less controllable. Therefore, it is necessary to label the elevation for multi-view videos. We used 100K 3D data with ground truth elevation to train our elevation estimator. Similar to the filtering process, we used DINO to extract features from 4 uniform views and averaged their features. A 2-layer MLP was trained on top of this as an elevation predictor. We then use it to label the elevation of all generated multi-view videos.
@@ 4.1 Implementation Details
Our VFusion3D model is trained on 128 NVIDIA A100 (80G) GPUs, and the training process takes approximately 6 days to complete. We use a total batch size of 1024, with 4 multi-view images in $128\times 128$ resolution used for supervision per batch, resulting in a total of 4096 multi-view images. The model is trained for 30 epochs with an initial learning rate of 4 $\times$ $10^{-4}$, following a cosine annealing schedule with a restart after first 10 epochs. The training begins with a warm-up of 3000 iterations, and the optimizer used is AdamW [33]. More details are presented in the Appendix 0.A.
@@ 4.2 Results and Comparisons
For text-to-3D generation, we employ EMU to generate images from text prompts, supporting text-image-3D reconstruction for our method. We use a common test set for evaluation, which contains 40 prompts from MVDream [49]. We compare VFusion3D with DreamFusion [41], Magic3D [26], ProlificDreamer [59], MVDream [49], OpenLRM [15, 16], and LGM [53]. For text-to-3D models, we use rendered videos from the MVDream website for evaluation. Results are presented in Table 2, and visual samples are shown in Figure 6. When considering feed-forward models, LGM demonstrates better text alignment, while VFusion3D exhibits stronger image alignment. Qualitatively, VFusion3D shows superior results compared to other feed-forward models, especially in terms of both color and shape 3D consistency.
@@ 4.3 Ablation Study
We conduct an ablation study on several design choices, which include: (1) the specifications for fine-tuning the EMU Video model, (2) the training strategies proposed in VFusion3D, and (3) the number of 3D data for fine-tuning pre-trained VFusion3D. We use two evaluation sets for our study. First, we report SSIM and LPIPS for 3D data evaluation, along with CLIP text similarity score and CLIP image similarity score for text-to-3D evaluation. The 3D data evaluations were conducted on a test dataset comprising 500 unseen 3D models. From different viewing angles, we use one image as input, render 32 novel views, and compare them against their respective ground truths. The text-to-3D evaluation follows the setting of text-to-3D experiments as in section 4.2.
@@ 5 Conclusion
This paper leverages a video diffusion model as a multi-view data generator, thereby facilitating the learning of scalable 3D generative models. We have showcased the potential of video diffusion models to function as a multi-view data engine, capable of generating an even infinite scale of synthetic data to support scalable training. Our proposed model, VFusion3D, which learns from synthetic data, has shown superior performance in the generation of 3D assets. Beyond its current state, VFusion3D is highly scalable and can scale with both the number of synthetic data and 3D data, paving new paths for 3D generative models. We hope our study can offer a potential pathway and valuable insights for the development of foundation 3D generative models.
Appendix 0.A Training Details
@@
Following the EMU Video [13], we freeze the spatial convolutional and attention layers of Emu Video, while only fine-tuning the temporal layers. We use the standard diffusion loss for this fine-tuning process. The Emu Video model is fine-tuned over a period of 5 days using 80 A100 GPUs, with a total batch size of 240 and a learning rate of 1 $\times$ $10^{-5}$. Although the 3D consistency continues to improve with extended fine-tuning, we do not observe any decline in visual quality. One possible explanation is the static nature of the spatial layers and the image-conditioned network, which ensures that the generated ${}^{\circ}$ videos maintain high fidelity with the high-frequency texture components of the input.
@@
The architecture of VFusion3D is identical to LRM [16]. In addition to the training details provided in the main paper, we use 0.95 as the second beta parameter of the AdamW optimizer [33]. We apply a gradient clipping of 1.0 and a weight decay of 0.05. This weight decay is only applied to weights that are neither bias nor part of normalization layers. For mixed precision training, we use Bfloat16 precision.
@@
Appendix 0.B Visualizations and Test-time Processing
@@
Following standard procedures [16, 53], we utilize a heuristic function to process all input images during testing. The initial steps involve eliminating the image background using rembg with isnet-general-use [43], then extracting the salient object. Following this, we adjust the size of the salient object to an appropriate scale and position it in the center of the input image. For rescaling, a smaller border ratio generally yields better visual quality in the generated results, whereas a larger border ratio tends to maintain higher fidelity to the original input image.
Appendix 0.C Failure Cases
@@
Appendix 0.D Conversation to Meshes

### v2

@@ Abstract
This paper presents a novel method for building scalable 3D generative models utilizing pre-trained video diffusion models. The primary obstacle in developing foundation 3D generative models is the limited availability of 3D data. Unlike images, texts, or videos, 3D data are not readily accessible and are difficult to acquire. This results in a significant disparity in scale compared to the vast quantities of other types of data. To address this issue, we propose using a video diffusion model, trained with extensive volumes of text, images, and videos, as a knowledge source for 3D data. By unlocking its multi-view generative capabilities through fine-tuning, we generate a large-scale synthetic multi-view dataset to train a feed-forward 3D generative model. The proposed model, VFusion3D, trained on nearly 3M synthetic multi-view data, can generate a 3D asset from a single image in seconds and achieves superior performance when compared to current SOTA feed-forward 3D generative models, with users preferring our results over $90\%$ of the time.
@@ 3.2 Video Diffusion Models as 3D Multi-view Data Generator
The generated multi-view videos follow a similar camera trajectory to cover a 360 azimuth range, but the elevation depends on the generated condition image, which is less controllable. Therefore, it is necessary to label the elevation for multi-view videos. We used 100K 3D data with ground truth elevation to train our elevation estimator. Similar to the filtering process, we used DINO to extract features from 4 uniform views and averaged their features. A 2-layer MLP was trained on top of this as an elevation predictor. We then use it to label the elevation of all generated multi-view videos.
@@ 4.1 Implementation Details
Our VFusion3D model is trained on 128 NVIDIA A100 (80G) GPUs, and the training process takes approximately 6 days to complete. We use a total batch size of 1024, with 4 multi-view images at a resolution of $128\times 128$ used for supervision per batch, resulting in a total of 4096 multi-view images. The input image has a resolution of 512 $\times$ 512. The model is trained for 30 epochs with an initial learning rate of 4 $\times$ $10^{-4}$, following a cosine annealing schedule with a restart after first 10 epochs. The training begins with a warm-up of 3000 iterations, and the optimizer used is AdamW [35]. More details are presented in the Appendix 0.B.
@@ 4.2 Results and Comparisons
For text-to-3D generation, we employ EMU to generate images from text prompts, supporting text-image-3D reconstruction for our method. We use a common test set for evaluation, which contains 40 prompts from MVDream [52]. We compare VFusion3D with DreamFusion [44], Magic3D [28], ProlificDreamer [65], MVDream [52], OpenLRM [16, 17], and LGM [56]. For text-to-3D models, we use rendered videos from the MVDream website for evaluation. Results are presented in Table 2, and visual samples are shown in Figure 6. When considering feed-forward models, VFusion3D exhibits stronger image and text alignment. Qualitatively, VFusion3D shows superior results compared to other feed-forward models, especially in terms of both color and 3D shape consistency.
@@ 4.3 Ablation Study
We conduct an ablation study on several design choices, which include: (1) the specifications for fine-tuning the EMU Video model, (2) the training strategies proposed in VFusion3D, and (3) the number of 3D data for fine-tuning pre-trained VFusion3D. We use two evaluation sets for our study. First, we report SSIM and LPIPS for 3D data evaluation, along with CLIP text similarity score and CLIP image similarity score for text-to-3D evaluation. The 3D data evaluations were conducted on a test dataset comprising 500 unseen 3D models. From different viewing angles, we use one image as input, render 32 novel views, and compare them against their respective ground truths. The text-to-3D evaluation follows the setting of text-to-3D experiments as in section 4.2. The VFusion3D model used in the ablation study 4.3 and analysis 4.4 is a variant that processes input images with a resolution of 224 $\times$ 224.
@@ 5 Conclusion
This paper leverages a video diffusion model as a multi-view data generator, thereby facilitating the learning of scalable 3D generative models. We have showcased the potential of video diffusion models to function as a multi-view data engine, capable of generating an even infinite scale of synthetic data to support scalable training. Our proposed model, VFusion3D, which learns from synthetic data, has shown superior performance in the generation of 3D assets. Beyond its current state, VFusion3D is highly scalable and can scale with both the number of synthetic data and 3D data, paving new paths for 3D generative models.
Acknowledgements
Junlin Han is supported by Meta. We would like to thank Jianyuan Wang, Luke Melas-Kyriazi, Yawar Siddiqui, Quankai Gao, Yanir Kleiman, Roman Shapovalov, Natalia Neverova, and Andrea Vedaldi for the insightful discussions and invaluable support.
Appendix 0.A arXiv Update Notes
This arXiv v2 version is the ECCV 2024 camera-ready version. Significant updates have been made compared to the arXiv v1 version, including: (1) Updating all major results of VFusion3D, which now uses a resolution of 512x512 for input images during both training and inference, compared to the 224$\times$224 resolution used in v1. (2) Incorporating benchmarking results on GSO [12]. (3) Adding experiments on SfM for fine-tuned Emu Video generated multi-view sequences. (4) Adding discussions on potential solutions for our current limitations with updated references.
Appendix 0.B Training Details
@@
Following the EMU Video [14], we freeze the spatial convolutional and attention layers of Emu Video, while only fine-tuning the temporal layers. We use the standard diffusion loss for this fine-tuning process. The Emu Video model is fine-tuned over a period of 5 days using 80 A100 GPUs, with a total batch size of 240 and a learning rate of 1 $\times$ $10^{-5}$. Although the 3D consistency continues to improve with extended fine-tuning, we do not observe any decline in visual quality. One possible explanation is the static nature of the spatial layers and the image-conditioned network, which ensures that the generated 360 videos maintain high fidelity with the high-frequency texture components of the input.
@@
The architecture of VFusion3D is identical to that of LRM [17], except for two minor modifications. The first is the use of DINOv2 [41] instead of DINO [4] to enhance image feature extraction capabilities. The second modification is that we have reduced the number of NeRF MLP layers to 4 instead of 10, as empirically suggested by the LRM ablation study.
In addition to the training details provided in the main paper, we use 0.95 as the second beta parameter of the AdamW optimizer [35]. We apply a gradient clipping of 1.0 and a weight decay of 0.05. This weight decay is only applied to weights that are neither bias nor part of normalization layers. We use Bfloat16 precision for training and inference.
@@
Appendix 0.C Visualizations and Test-time Processing
@@
Following standard procedures [17, 56], we utilize a heuristic function to process all input images during testing. The initial steps involve eliminating the image background using rembg with isnet-general-use [46], then extracting the salient object. Following this, we adjust the size of the salient object to an appropriate scale and position it in the center of the input image.
Appendix 0.D Benchmarking on Public Datasets
We present a comparison using the GSO [12] dataset, from which we render 996 objects (excluding some redundant or very similar objects) with each object having 16 rendered views. These views are rendered with a fixed elevation angle of 20. The azimuth angle is sampled uniformly between 0 and 360. The input image is the first rendered view, namely, the view with an azimuth of 0.
Results are shown in Table 7, which demonstrates that VFusion3D significantly outperforms the baselines by large margins. This further demonstrates that utilizing synthetic data to scale 3D generative models holds great promise.
[…]
OpenLRM 14.21 0.820 0.226
LGM 13.53 0.811 0.254
VFusion3D 20.89 0.848 0.127
Table 7: Performance Comparison on the GSO Dataset: VFusion3D significantly outperforms the baselines by substantial margins.
Appendix 0.E Discussions on Limitations and Potential Solutions
@@
One way to assess the failure rate and check the pixel-level multi-view correspondence is to run Structure from Motion (SfM) algorithms, which require pixel-level correspondences, on generated multi-view frames. We run VGGSfM [62] on 1000 fine-tuned EMU Video-generated multi-view sequences. Among these sequences, VGGSfM is able to obtain reasonable pose estimations (at least accurate poses for 8 views) over 70% of the time. Examples are presented in Figure 9. This demonstrates that the generated sequences generally have reasonable pixel-level 3D correspondence, but there is significant room for further improvements. Enforcing stronger multi-view consistency constraints [61, 20, 69, 58] during the video fine-tuning stage is a plausible solution. This could involve the use of epipolar constraints and 3D-aware Plücker ray embeddings.
@@
Figure 9: VGGSfM reconstruction on all 16 generated frames. Our fine-tuned EMU Video typically produces multi-view sequences with reasonable pixel-level correspondences.
Appendix 0.F Conversation to Meshes